# High-Capacity and Low-PAPR BICM-OFDM Systems Using Non-Equiprobable and Non-Uniform Constellation Shaping With Clipping and Filtering

### v1

@@ III-A Total BMI
where $N_{0}$ represents the power spectral density of AWGN, which is related to $\gamma$ by $\gamma=E_{s}/N_{0}$.
III-B Parameter Optimization
For a given pair of modulation order $M$ and operating SNR $\gamma$, the optimization problem in our NENU constellation is formulated as
@@ III-B Parameter Optimization
By evaluating the BMI over all parameter pairs of $(N,\rho)$, the optimal constellation can be identified. (In practice, we may apply the quantized value of $\rho$ with the step size of $0.01$.) Note that when the channel is modeled as AWGN, the expectation over the Gaussian noise in the BMI calculation (8) can be efficiently computed using Gaussian–Hermite quadrature (GHQ) [35].
III-C Loss From the Shannon Limit
Figure 4: Loss in achievable information rate from the Shannon limit for the BICM systems using the proposed constellations with various modulation orders.
The gain achieved by constellation shaping is typically assessed by the difference between the achievable information rate and the Shannon capacity, known as the loss from the Shannon limit [36]. When the two identical $M$-ary PAM constellations form $M^{2}$-ary QAM, the loss from the Shannon limit is defined as a function of SNR $\gamma$ by
@@ III-C Loss From the Shannon Limit
The loss from the Shannon limit of the proposed shaping scheme is shown in Fig. 4 for various modulation orders ranging from $M^{2}=16$ to $M^{2}=1\,048\,576$. Each plotted point is obtained by selecting the optimal parameter set $\left(N,\rho\right)$ at a given SNR. As a comparison, the minimum loss achieved by the uniform $M^{2}$-QAM with the optimized modulation order is indicated as a gray dashed line. At low SNRs, the performance of the proposed shaping scheme coincides with that of conventional QAM, due to the optimality of low-order constellations such as QPSK and uniform 16-QAM in this SNR regime. However, as the SNR increases, the gap between the BMI of standard QAM and the Shannon limit becomes noticeable. In contrast, the proposed NENU constellation consistently benefits from higher modulation orders across all SNRs. This indicates that, due to its flexible control over the number of signal points, properly designed NENU constellations successfully avoid excessive geometric density that would otherwise degrade the BMI. In fact, the proposed NENU constellation enhances the BMI because of an effect similar to probabilistic shaping.
@@ III-D Comparison With Other Geometric Shaping Schemes
We next compare the proposed NENU constellations with other constellations. For this purpose, we fix the modulation order as $M^{2}=1024$ in what follows. We consider the following numerically optimized constellations for comparison:
@@
  • The two-dimensional (2D) $1024$-NUC developed in [7].
  • The uniform $M^{2}$-ary QAM with $M^{2}\in\{64,256,1024\}$
  • The one-dimensional shaping without constellation merging, i.e., ${\cal A}\left(M,M,\rho\right)$ with $M^{2}\in\{64,256,1024\}$ (indicated as shaped QAM).
  • The proposed one-dimensional NENU constellation ${\cal A}\left(32,N,\rho\right)$.
The results are compared in Fig. 5. We first observe that the optimal modulation order in BICM is contingent upon the channel SNR. The shaped QAM without constellation merging has been confirmed to demonstrate improved performance compared to uniform QAM [23]. However, due to the constraint of constellation sizes to powers of two, these schemes exhibit significant gaps from the Shannon limit at specific SNR regions, particularly around 15 dB and 20 dB. In the case of the proposed shaping scheme, these gaps can be maintained at a minimal level across a wide range of SNR. This suggests that employing shaping with a flexible constellation size is advantageous for rate-adaptive BICM design. The gain achieved through the proposed shaping scheme is comparable to that of the one-dimensional NUC in ATSC 3.0 across all the SNRs, even though the latter is obtained by a more complex optimization process.
When comparing the results with the two-dimensional constellation of [7], performance gaps of approximately $0.08$ bits per symbol are observed, highlighting the limitations inherent in one-dimensional constellation design. However, one-dimensional modulation schemes offer practical advantages in very high data-rate regimes, since the demapping process for two-dimensional constellations requires a considerable increase in computational cost with higher modulation orders, as is discussed in the next subsection.
@@ III-E Demapping Complexity
For efficient decoding, it is necessary to compute bit log-likelihood ratio (LLR) and this dominates the computational complexity at the receiver especially when irregular constellations are adopted. In what follows, we discuss the demapping complexity for general QAM constellations  [^3] There have been several attempts to simplify the soft-output demappers for both one-dimensional and two-dimensional non-uniform constellations [37, 38, 39, 40], but they rely on specific geometric structures of constellations. .
Let us assume that the transmitted symbols are selected from a set of specific complex-valued (i.e., two-dimensional) constellation set $\mathcal{A}$, with the number of points represented by $|\mathcal{A}|$. The LLR value for the $i$th bit level is computed as
$L_{i}$ $=\log\left[\frac{\sum_{x\in\mathcal{A}_{0}^{(i)}}\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}{\sum_{x\in\mathcal{A}_{1}^{(i)}}\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}\right]$
$\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{A}_{1}^{\left(i\right)}}\left\{|y-x|^{2}\right\}-\min_{x\in\mathcal{A}_{0}^{\left(i\right)}}\left\{|y-x|^{2}\right\}\right),$ (12)
where $y\in{\mathbb{C}}$ denotes the received complex-valued symbol. We focus only on the number of required multiplication operations as the complexity associated with comparison and addition operations is considered negligible in comparison to that of multiplication operations.
On the other hand, if $\mathcal{A}$ is a set of one-dimensional constellation points, the LLR value can be expressed as
where $y\in{\mathbb{R}}$ is the received real-valued symbol and $x$ is chosen from one-dimensional (i.e., real-valued) constellation.
In the case of a two-dimensional constellation, from (12) it is necessary to compute the squared Euclidean distances between the received symbol and each of the $|\mathcal{A}|$ constellation points. This computation entails performing $2|\mathcal{A}|$ real multiplications. Furthermore, the cardinality of the two-dimensional constellations is $|\mathcal{A}|=M^{2}$. On the other hand, in the case of a one-dimensional constellation, calculating (13) requires $|\mathcal{A}|$ real multiplications, and the corresponding cardinality is $|\mathcal{A}|=M$. Therefore, as $M$ increases for higher spectral efficiency, the use of a one-dimensional constellation should be critical in view of practical implementation. This reasoning may elucidate the rationale behind the adoption of one-dimensional constellations for high-order scenarios in practical wireless communication systems such as ATSC 3.0, despite its lower achievable gain.
@@
The proposed NENU constellation also offers a more efficient demapping implementation due to its reduced constellation size from $M$, since the demapping complexity is proportional to the actual number of constellation points $N$. With the modulation order set as $M=32$, the optimal constellation size $N$ and the design parameter $\rho$ as a function of channel SNR $\gamma$ are illustrated in Fig. 6. It is observed that the required number of distinct signal points decreases as the SNR decreases. It thus demonstrates that the proposed NENU constellation not only achieves near-optimal gain but also minimizes the demapping complexity.
@@ IV Application to BICM-OFDM With CAF and CNC
The design of wireless signals with low PAPR continues to be a prominent and challenging field of research [41, 42, 43, 44], and a major practical drawback of constellation shaping is its increasing PAPR as it targets Gaussian distribution. On the other hand, BICM-OFDM is a de facto standard for high data rate wireless communications over frequency-selective fading channels, but due to its multicarrier nature, its signal has high PAPR regardless of the constellations employed. Thus, PAPR reduction should be necessary for better amplifier efficiency. Among many PAPR reduction approaches developed for OFDM, we suggest employing CAF [13], as it is directly applicable to the standard OFDM systems. The major drawback is its nonlinear distortion, which should be compensated for at the receiver. Therefore, we also suggest using CNC [14], which allows us to take advantage of one-dimensional constellations for efficient implementation [15]. We demonstrate that the proposed NENU constellations can be flexibly optimized for such systems, achieving notable gains under practical scenarios.
@@
${\cal Z}_{M,N,\rho}\triangleq\{X_{I}+jX_{Q}\,|\,X_{I},X_{Q}\in{\cal A}\left(M,N,\rho\right)\}.$
@@ IV-A OFDM Transmitter With CAF
We follow the description of [13] for CAF and that of [15] for CNC with an extension to coded OFDM. The block diagram of the OFDM transmitter with CAF is illustrated in Fig. 7. The information sequence is first encoded by a binary channel code with code rate $R_{c}\in(0,1]$. The codeword is bit-wise interleaved and then mapped onto the corresponding NENU constellations. Let $\mathbf{Z}=\left(Z_{0},Z_{1},\ldots,Z_{N_{c}-1}\right)\in\mathcal{Z}_{M,N,\rho}^{N_{c}}$ denote a vector of QAM symbols carried by each $N_{c}$-subcarrier OFDM signal, where $Z_{k}=X_{I,k}+jX_{Q,k}$ with $X_{I,k}$ and $X_{Q,k}$ representing NENU PAM symbol on the $k$th subcarrier. The $J$-times oversampled OFDM symbol $\mathbf{S}=(S_{0},S_{1},...,S_{JN_{c}-1})\in{\mathbb{C}}^{JN_{c}}$ is generated by performing $JN_{c}$-point inverse discrete Fourier transform (IDFT) on the zero-padded symbol vector $\mathbf{Z}^{\prime}=(\mathbf{Z},\underbrace{0,0,\ldots,0}_{(J-1)N_{c}})\in{\mathbb{C}}^{JN_{c}}$ as follows:444 In practice, DFT and IDFT are implemented using the fast Fourier transform (FFT) algorithm by setting the number of points $JN_{c}$ as the powers of two.
$S_{n}=\frac{1}{\sqrt{JN_{c}}}\sum_{k=0}^{JN_{c}-1}Z^{\prime}_{k}e^{j2\pi k\frac{n}{JN_{c}}},\quad n=0,1,\ldots,JN_{c}-1.$ (14)
To restrict the peak power, these time domain signal samples are clipped by soft-envelope limiter [13] as
$\tilde{S}_{n}=\begin{cases}S_{n},&\text{ for }\left|S_{n}\right|\leq A_{\text{max}},\\ A_{\text{max}}e^{j\arg S_{n}},&\text{ for }\left|S_{n}\right|>A_{\text{max}},\end{cases}$ (15)
where $A_{\text{max}}$ represents the maximum envelope level. We define the clipping ratio $\gamma_{\text{CR}}$ as
$\gamma_{\text{CR}}\triangleq\frac{A_{\text{max}}}{\sqrt{P_{\text{in}}}},$ (16)
where $P_{\mathrm{in}}\triangleq\mathbb{E}\left[\|\mathbf{Z}\|^{2}\right]/N_{c}$ is the average power of the OFDM signal prior to CAF. We then perform $JN_{c}$-point discrete Fourier transform (DFT) on the clipped signals to transform them back to the frequency-domain symbols represented by $\mathbf{U}^{\prime}=\left(U^{\prime}_{0},U^{\prime}_{1},\ldots,U^{\prime}_{JN_{c}-1}\right)\in{\mathbb{C}}^{JN_{c}}$, where

### v2

@@ III-A Bit-wise Mutual Information
with $N_{0}$ representing the power spectral density of the AWGN, which is related to $\gamma$ by $\gamma=E_{s}/N_{0}$. Note that The above integral can be efficiently evaluated using Gaussian–Hermite quadrature (GHQ) [alvarado2018achievable]. Accordingly, the achievable information rate of a BICM system is defined as the sum of the BMIs over all binary channels:
$C\left(\mathcal{A},\gamma\right)=\sum_{i=0}^{m-1}C_{i}(\mathcal{A},\gamma),$ (11)
measured in bits per dimension.
TABLE II: BMI of individual binary channels for $(32,N)$-PAM with $\rho=1$. The bit $b^{(0)}$ corresponds to the MSB.
NN (SNR) $b^{(0)}$ $b^{(1)}$ $b^{(2)}$ $b^{(3)}$ $b^{(4)}$ Total
32 (10 dB) 0.6940 0.5163 0.2291 0.0741 0.0268 1.5403
10 (10 dB) 0.7518 0.5915 0.1739 0.0695 0.0262 1.6128
32 (20 dB) 0.9013 0.8453 0.7193 0.4800 0.2239 3.1699
10 (20 dB) 0.9990 0.9953 0.4812 0.2241 0.1158 2.8154
[…]
III-B Effect of Constellation Point Merging on BMI
To elucidate the validity of the proposed merging operation, we first demonstrate its impact on the BMI of individual binary channels. As an example of $M=32$ (corresponding to five binary channels), Table II lists the BMI values of individual binary channels for $(32,32)$-PAM and $(32,10)$-PAM ($N=32$ and $N=10$), each evaluated at SNRs of $10$ dB and $20$ dB, where $b^{(0)}$ and $b^{(4)}$ correspond to the MSB and the LSB, respectively. It is observed in both SNR cases that the BMIs of the lower-order bits, such as $b^{(3)}$ and $b^{(4)}$, are reduced by the merging process, while those of the higher-order bits, $b^{(0)}$ and $b^{(1)}$, are increased. Therefore, the proposed merging operation enhances the reliability of the higher-order bits at the expense of that of the lower-order bits. In the case of lower SNR ($10$ dB), in particular, the lower-order bits are inherently unreliable and contribute only marginally to the total BMI. As a result, the loss in their BMIs due to merging is relatively small, whereas the enhanced separability among constellation points successfully increases the BMIs of the higher-order bits, leading to an overall improvement in the total BMI.
III-C Loss From the Shannon Limit
Figure 4: Loss in achievable information rate from the Shannon limit for the BICM systems using the proposed NENU constellations with various modulation orders.
We evaluate the gain achieved by the proposed shaping scheme when its parameters are properly optimized. For a given pair of modulation order $M$ and operating SNR $\gamma$, the optimization problem for our NENU constellation is formulated as
@@ III-C Loss From the Shannon Limit
In this work, constellations with varying numbers of distinct points $N\in\{2,4,\ldots,M\}$ are generated for each quantized value of $\rho$, using a step size of $0.01$, in accordance with the proposed merging principle. The optimal constellation is determined by evaluating the BMI in (11) across all candidate parameter pairs $(N,\rho)$. It is important to note that if $\rho$ is fixed for a specific target SNR and $N$ is subsequently adjusted to enhance the BMI, the optimization process would be simplified. However, this approach may lead to a suboptimal constellation.
The gain achieved by constellation shaping is typically assessed by the difference between the achievable information rate and the Shannon limit, known as the loss from the Shannon limit [montorsi2016design]. When the two identical $M$-ary PAM constellations form $M^{2}$-ary QAM, the loss from the Shannon limit is defined as a function of SNR $\gamma$ by
@@
The loss from the Shannon limit of the proposed shaping scheme is shown in Fig. 4 for various modulation orders ranging from $M^{2}=16$ to $M^{2}=1\,048\,576$. Each plotted point is obtained by selecting the optimal parameter set $\left(N,\rho\right)$ at a given SNR. As a comparison, the minimum loss achieved by the uniform $M^{2}$-QAM with the optimal modulation order is indicated as a gray dashed line. At low SNRs, the performance of the proposed shaping scheme coincides with that of conventional QAM, due to the optimality of low-order constellations such as QPSK and standard uniform 16-QAM in this SNR regime. However, as the SNR increases, the gap between the BMI of standard QAM and the Shannon limit becomes increasingly pronounced. On the other hand, the proposed NENU constellation achieves a minimal loss from the Shannon limit, particularly in the high-SNR regimes, by exploiting higher modulation orders. Specifically, losses of less than $0.1$ bits per symbol are observed for modulation orders of 16 38416\,384 and above. These remarkable gains are attributed to the high-resolution Gaussian approximation enabled by the huge number of constellation points. Moreover, higher-order NENU constellations consistently outperform lower-order ones across the entire SNR range. This confirms that the proposed merging scheme adaptively reduces the effective constellation size of high-order NENU constellations in the low-SNR regime.
@@ III-D Comparison With Other Geometric Shaping Schemes
We next compare the proposed NENU constellations with other representative modulation schemes. The following constellations are considered for comparison:
  • The uniform $M^{2}$-ary QAM with $M^{2}\in\{64,256,1024\}$
  • The proposed one-dimensional shaping without constellation point merging, i.e., ${\cal A}\left(M,M,\rho\right)$ with $M^{2}\in\{64,256,1024\}$ (indicated as shaped QAM).
  • The proposed NENU constellation ${\cal A}\left(32,N,\rho\right)$.
@@
  • The two-dimensional (2D) $1024$-NUC developed in [sillekens2022high].
The corresponding loss from the Shannon limit is compared in Fig. 5, where all shaping schemes other than standard uniform QAM are optimized at each plotted SNR according to their respective design or optimization algorithms.
We begin with a brief discussion of the design complexity in terms of the number of real-valued variables to be optimized (i.e., the degrees of freedom). Let $m^{\prime}=2m=\log_{2}M^{2}$ denote the number of binary channels (i.e., the maximum spectral efficiency in bits per symbol). For the one-dimensional NUC in ATSC, $2^{m^{\prime}/2-1}=2^{m-1}$ real-valued PAM amplitudes need to be optimized under an origin-symmetry constraint. On the other hand, the two-dimensional NUC in [sillekens2022high] is designed directly in the complex plane under a quadrant symmetry. In this case, $2^{m^{\prime}-2}=2^{2(m-1)}$ complex-valued constellation points in a single quadrant must be optimized, which corresponds to $2^{2m-1}$ real-valued variables for the I/Q components. In contrast, for the proposed NENU constellation, all PAM amplitudes are optimized using only two real-valued parameters irrespective of $m$, highlighting its scalability for high-spectral-efficiency regimes.
As shown in Fig. 5, the analysis of the achievable rate reveals that the optimal modulation order in BICM is influenced by the channel SNR. The implementation of shaped QAM without merging constellation points has been validated to provide better performance compared to standard uniform QAM [kurihara2024design]. However, the limitation of constellation sizes being restricted to powers of two results in notable gaps from the Shannon limit at certain SNR ranges, specifically around 15 dB and 20 dB. In contrast, the proposed shaping scheme maintains these gaps at a minimal level over a broad range of SNRs. This indicates that using shaping with the ability to flexibly adjust both the constellation size and its distribution is beneficial for a rate-adaptive BICM design. The performance gain from the proposed shaping scheme is comparable to that of the one-dimensional NUC in ATSC 3.0 across all SNR ranges, despite the latter being achieved through a more complex optimization process for a specific target SNR or information rate.
When comparing the results with the two-dimensional constellation of [sillekens2022high], performance gaps of approximately $0.08$ bits per symbol are observed, which highlights the inherent limitations of one-dimensional constellation design. However, one-dimensional modulation schemes offer practical advantages in very high data-rate regimes, since the demapping process for two-dimensional constellations requires a substantial increase in computational cost with higher modulation orders, as is discussed in the next subsection.
@@ III-E Demapping Complexity
In BICM systems, soft demapping is carried out by computing the bit log-likelihood ratios (LLRs) for each binary channel, which dominates the computational complexity at the receiver. The required computational effort strongly depends on the geometric structure of the underlying constellation and becomes increasingly critical as the modulation order increases. Accordingly, we analyze the LLR computation complexity for general one-dimensional and two-dimensional constellations and then discuss the corresponding complexity of the proposed NENU constellation.
Let us assume that the transmitted symbols are selected from a set of specific complex-valued (i.e., two-dimensional) constellation $\mathcal{X}$, with the number of points represented by $|\mathcal{X}|$. The LLR value for the $i$th binary channel is computed as
$L_{i}$ $=\log\left[\frac{\sum_{x\in\mathcal{X}_{0}^{(i)}}p(x)\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}{\sum_{x\in\mathcal{X}_{1}^{(i)}}p(x)\exp\left(-\frac{|y-x|^{2}}{N_{0}}\right)}\right]$
$\approx\frac{1}{N_{0}}\left(\min_{x\in\mathcal{X}_{1}^{\left(i\right)}}\left\{|y-x|^{2}\right\}-\min_{x\in\mathcal{X}_{0}^{\left(i\right)}}\left\{|y-x|^{2}\right\}\right),$ (14)
where $y\in{\mathbb{C}}$ denotes the received complex-valued symbol and $\mathcal{X}_{b}^{(i)}\subset\mathcal{X}$ represents the set of constellation points in the $i$th binary channel with the bit specified by $b\in\mathbb{F}_{2}$. On the other hand, for a one-dimensional constellation, where $\mathcal{X}$ is composed of two identical and orthogonal PAM constellations $\mathcal{A}$, the LLR value can be expressed as
where $y\in{\mathbb{R}}$ is the received real-valued symbol, $\mathcal{A}_{b}^{(i)}\subset\mathcal{A}$ denotes the set of constellation points in the $i$th binary channel with the bit specified by $b\in\mathbb{F}_{2}$, and $x$ is thus chosen from a one-dimensional (i.e., real-valued) constellation.
We focus only on the number of required multiplication operations, since the complexity associated with comparison and addition operations is negligible compared to that of multiplication operations. In the case of a two-dimensional constellation, from (14) it is necessary to compute the squared Euclidean distances between the received symbol and each of the $|\mathcal{X}|=M^{2}$ candidate constellation points. This computation entails performing $2|\mathcal{X}|$ real multiplications. Meanwhile, in the case of a one-dimensional constellation, calculating (15) requires $|\mathcal{A}|$ real multiplications. The square root of the cardinality of a two-dimensional constellation, $|\mathcal{A}|=M$, represents the corresponding cardinality in this case. Therefore, as $M$ increases to support higher spectral efficiency, the use of a one-dimensional constellation should be critical from an implementation complexity perspective. This reasoning may elucidate the rationale behind the adoption of one-dimensional constellations for high-order scenarios in practical wireless communication systems such as ATSC 3.0.
@@
The proposed NENU constellation also offers a more efficient demapping implementation due to its reduced constellation cardinality. In the LLR computation in (15), since Euclidean distances are compared among $|\mathcal{A}(M,N,\rho)|=N$ candidate constellation points, the number of required real multiplications is given by $N$. With the modulation order set to $M=32$, the optimal constellation size $N$ and the design parameter $\rho$ as a function of channel SNR $\gamma$ are illustrated in Fig. 6. It is observed that the required number of distinct signal points reduces as the SNR decreases. It thus demonstrates that the proposed NENU constellation not only achieves near-optimal gain but also minimizes the demapping complexity.
We note that several simplified demapping algorithms have been developed for both one-dimensional and two-dimensional NUCs [fuentes2015low, hong2020backward, hong2022enhanced, barrueco2017condensation, duan2021non, barrueco2018low]. Quadrant search reduction (QSR) [fuentes2015low] reduces the number of the required Euclidean distance computations by discarding many symbols in different quadrants that have negligible impact on the LLR. Virtual point searching (VPS) [hong2020backward, hong2022enhanced], on the other hand, approximates a general two-dimensional constellation by using virtual points located on several radial axes, similar to the amplitude-phase shift keying (APSK) constellation. While this approximation can significantly reduce the demapping complexity, it could lead to a substantial degradation in the achievable information rate. Condensed symbol reduction (CSR) [barrueco2017condensation, duan2021non, barrueco2018low] reduces the demapping complexity by exploiting the clustering of constellation points that typically appears in NUCs optimized for low code rates, where each cluster is treated as a single symbol in the LLR computation. CSR can be applied to both one-dimensional and two-dimensional NUCs. In the case of two-dimensional NUCs, it is sometimes combined with VPS or QSR to further reduce the complexity [fuentes2015low, hong2020backward, hong2022enhanced].
Nevertheless, the most effective reduction in demapping complexity is still achieved by exploiting I/Q independence, which naturally favors one-dimensional constellations especially for high modulation orders. In this context, a one-dimensional NUC with a carefully designed CSR-based demapper may achieve performance comparable to the proposed NENU constellation. However, it should rely on heuristic procedures that determine the initial constellation points and perform the clustering process. In contrast, since the proposed NENU constellation is uniquely described by a simple mathematical expression of (5), it ensures systematic reproducibility and enables a significantly simpler optimization process.
@@ IV Application to BICM-OFDM With CAF and CNC
The design of wireless signals with low PAPR continues to be a prominent and challenging field of research [9725261, 9747919, 9962325, 10530524]. BICM-OFDM is a de facto standard for high data rate wireless communications over frequency-selective fading channels, but due to its multicarrier nature, its signal has high PAPR regardless of the constellations employed. Thus, PAPR reduction should be necessary for better power amplifier efficiency. Among many PAPR reduction approaches developed for OFDM, we suggest employing CAF [ochiai2002performance], as it is directly applicable to the BICM-OFDM system using the proposed NENU constellation. The major drawback is its nonlinear distortion, which should be compensated for at the receiver. Therefore, we address the use of CNC [chen2003iterative, sun2021], which is also compatible with our constellation shaping. We demonstrate that the proposed NENU constellations can be flexibly optimized for such systems, achieving notable gains under practical scenarios.
@@
${\cal X}_{M,N,\rho}\triangleq\{A_{I}+jA_{Q}\,|\,A_{I},A_{Q}\in{\cal A}\left(M,N,\rho\right)\}.$
@@ IV-A OFDM Transmitter With CAF
We follow the description of [ochiai2002performance] for CAF and that of [sun2021] for CNC with an extension to coded OFDM. The block diagram of the OFDM transmitter with CAF is illustrated in Fig. 7. The information sequence is first encoded by a binary channel code with code rate $R_{c}\in(0,1]$. The codeword is bit-wise interleaved and then mapped onto the corresponding NENU constellations. Let $\mathbf{X}=\left(X_{0},X_{1},\ldots,X_{N_{c}-1}\right)\in\mathcal{X}_{M,N,\rho}^{N_{c}}$ denote a vector of QAM symbols carried by each $N_{c}$-subcarrier OFDM signal, where $X_{k}=A_{I,k}+jA_{Q,k}$ with $A_{I,k}$ and $A_{Q,k}$ representing NENU PAM symbol on the $k$th subcarrier. The $J$-times oversampled OFDM symbol $\mathbf{S}=(S_{0},S_{1},...,S_{JN_{c}-1})\in{\mathbb{C}}^{JN_{c}}$ is generated by performing $JN_{c}$-point inverse discrete Fourier transform (IDFT) on the zero-padded symbol vector $\mathbf{X}^{\prime}=(\mathbf{X},\underbrace{0,0,\ldots,0}_{(J-1)N_{c}})\in{\mathbb{C}}^{JN_{c}}$ as follows: [^4] In practice, DFT and IDFT are implemented using the fast Fourier transform (FFT) algorithm by setting the number of points $JN_{c}$ as the powers of two.
$S_{n}=\frac{1}{\sqrt{JN_{c}}}\sum_{k=0}^{JN_{c}-1}X^{\prime}_{k}e^{j2\pi k\frac{n}{JN_{c}}},\quad n=0,1,\ldots,JN_{c}-1.$ (16)
To restrict the peak power, these time domain signal samples are clipped by soft-envelope limiter [ochiai2002performance] as
$\tilde{S}_{n}=\begin{cases}S_{n},&\text{ for }\left|S_{n}\right|\leq r_{\text{max}},\\ r_{\text{max}}e^{j\arg S_{n}},&\text{ for }\left|S_{n}\right|>r_{\text{max}},\end{cases}$ (17)
where $r_{\text{max}}$ represents the maximum envelope level. We define the clipping ratio $\gamma_{\text{CR}}$ as
$\gamma_{\text{CR}}\triangleq\frac{r_{\text{max}}}{\sqrt{P_{\text{in}}}},$ (18)
where $P_{\mathrm{in}}\triangleq\mathbb{E}\left[\|\mathbf{X}\|^{2}\right]/N_{c}$ is the average power of the OFDM signal prior to CAF. We then perform $JN_{c}$-point discrete Fourier transform (DFT) on the clipped signals to transform them back to the frequency-domain symbols represented by $\mathbf{U}^{\prime}=\left(U^{\prime}_{0},U^{\prime}_{1},\ldots,U^{\prime}_{JN_{c}-1}\right)\in{\mathbb{C}}^{JN_{c}}$, where